# TSPO: Temporal Sampling Policy Optimization for Long-form Video Language Understanding

### v2

@@ Introduction
Comparison with Keyframe Training Method. Both FrameVOYAGER (Yu et al. 2024) and MLLM-VFS (Hu et al. 2025b) are recent training-based methods that require offline keyframe ranking or labeling to supervise the keyframe selector. Compared with them, our TSPO offers two advantages: 1) RL modeling: we jointly model the keyframe selection and language generation from an RL perspective, enabling end-to-end optimization with off-the-shelf video QA data, without requiring additional frame annotations like MLLM-VFS. 2) Superior sampling strategy: FrameVOYAGER requires random pre-processing sampling from frame combinations, while our approach is on-policy, dynamically sampling based on the current policy, which progressively refines the selection strategy. For a fair comparison, since neither method has been open-source, we use their reported results from their papers and adapt our TSPO to the same settings. As shown in Tab. 3, TSPO achieves enhanced performance despite using less training data.
@@
Exploring SFT* for Keyframe. We investigate the possibility of end-to-end (E2E) optimization for the keyframe selector through SFT and compare it with our proposed TSPO. To enable gradient propagation from the LLM output to the temporal agent, we utilize the Gumbel-Softmax technique (Wei et al. 2023), which enables the selected vision token to have gradients. For SFT training, we randomly select 30K samples from LLaVA-Video-178K while keeping the MLLM parameters frozen. As shown in Tab. 5, our experimental results demonstrate that TSPO consistently outperforms SFT*, which indicates TSPO’s advantages, including its capacity to explore diverse sampling strategies and utilize more direct reward signals.
Inference Efficiency. In this study, we fix the candidate frame number $V_{c}$ to 128 (1FPS in our main setting), which avoids the effect of varying video lengths. As shown in Tab. 6, our efficiency can be demonstrated from the following aspects: 1) with the same 64 frames, ours achieves a consistent performance gain over the baseline; 2) with fewer yet informative frames, we maintain gains over baseline while requiring only half the number of tokens and reducing the LLM time to 50% of the original one; 3) for keyframe extraction time, our approach saves 90% of the time compared to CoS (Hu et al. 2025a), yet achieves performance gain. This demonstrates our efficiency in handling long videos.
Qualitative Results. Fig. 5 demonstrates the visualization results of keyframe selection and model responses. Our trained temporal agent is shown to achieve two capabilities: basic object recognition, e.g., “short-haired woman” or “airplane”, and temporal event relationship comprehension, e.g., “entering the museum” or “appeared before”. When the short-haired woman is localized, preceding contextual frames (the first scenes that the woman visits) are simultaneously captured, which confirms that our TSPO effectively guides the temporal agent to learn complex query-event correlation capacity. Furthermore, our precise keyframe selection enables the MLLM to generate accurate responses. This capability stems from our joint modeling of keyframe selection and language generation, coupled with end-to-end optimization via language supervision signals.
@@ Conclusion
This paper proposes a Temporal Sampling Policy Optimization framework, which addresses the unsupervised and non-differentiable challenge of sparse frame sampling in Video-MLLMs. Our method features two innovations: (1) an RL mechanism that models keyframe selection and language generation as a joint decision-making process, performing end-to-end optimization of sparse sampling; (2) a TSPO-targeted training data construction pipeline with dual-style data, incorporating the establishment of rule-based answering accuracy and temporal locating reward mechanisms. Extensive comparison experiments and ablation studies validate the effectiveness and generalizability of our method.

### v4

@@ Introduction
Comparison with Keyframe Training Method. Both FrameVOYAGER (Frame-Voyager) and MLLM-VFS (MLLM-VFS) are recent training-based methods that require offline keyframe ranking or labeling to supervise the keyframe selector. Compared with them, our TSPO offers two advantages: 1) RL modeling: we jointly model the keyframe selection and language generation from an RL perspective, enabling end-to-end optimization with off-the-shelf video QA data, without requiring additional frame annotations like MLLM-VFS. 2) Superior sampling strategy: FrameVOYAGER requires random pre-processing sampling from frame combinations, while our approach is on-policy, dynamically sampling based on the current policy, which progressively refines the selection strategy. For a fair comparison, since neither method has been open-source, we use their reported results from their papers and adapt our TSPO to the same settings. As shown in Tab. 3, TSPO achieves enhanced performance despite using less training data.
@@
Exploring SFT* for Keyframe. We investigate the possibility of end-to-end (E2E) optimization for the keyframe selector through SFT and compare it with TSPO. We utilize the Gumbel-Softmax technique (GRANSAC), which enables the selected vision token to have gradients (VidF4) and can be E2E trained. For SFT* training, we randomly select 30K samples from LLaVA-Video-178K while keeping the MLLM parameters frozen. As shown in Tab. 5, our experimental results demonstrate that TSPO consistently outperforms SFT*, which indicates TSPO’s advantages, including its capacity to explore diverse sampling strategies and utilize more direct reward signals.
Inference Efficiency. In this study, we fix the candidate frame number $V_{c}$ to 128 (1FPS in our main setting), which avoids the effect of varying video lengths. As shown in Tab. 6, our efficiency can be demonstrated from the following aspects: 1) with the same 64 frames, ours achieves a consistent performance gain over the baseline; 2) with fewer yet informative frames, we maintain gains over baseline while requiring only half the number of tokens and reducing the LLM time to 50% of the original one; 3) for keyframe extraction time, our approach saves 90% of the time compared to CoS (hu2025cos), yet achieves performance gain. This demonstrates our efficiency in handling long videos.
Qualitative Results. Fig. 5 demonstrates the visualization results of keyframe selection and model responses. Our trained temporal agent is shown to achieve two capabilities: basic object recognition, e.g., “short-haired woman” or “airplane”, and temporal event relationship comprehension, e.g., “entering the museum” or “appeared before”. When the short-haired woman is localized, preceding contextual frames (the first scenes that the woman visits) are simultaneously captured, which confirms that our TSPO effectively guides the temporal agent to learn complex query-event correlation capacity. Furthermore, our precise keyframe selection enables the MLLM to generate accurate responses.
@@ Conclusion
This paper proposes a Temporal Sampling Policy Optimization framework, which addresses the unsupervised and non-differentiable challenge of sparse frame sampling in Video-MLLMs. We propose an RL framework to optimize sparse frame sampling in an end-to-end manner, and propose a TSPO-targeted training data construction pipeline. Extensive comparison experiments and ablation studies validate the effectiveness and generalizability of our method.
Acknowledgments
This work was supported in part by National Science and Technology Major Project under Grant 2023ZD0121300, National Natural Science Foundation of China under Grants 62088102, U24A20325, 12326608, 62522102 and 62373043, and Key Research and Development Plan of Shaanxi Province under Grant 2024PT-ZCK-80.